OpenStudy (anonymous):

Find the exact area of the circle with the given radius. r = 3 mm A. 9 mm2 B. 9π mm2 C. 3π mm2 D. 9.42 mm2

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

whats the formula to finding the area of circles

OpenStudy (anonymous):

\(\pi r^2\) \(3^2\) What's 3 * 3?

OpenStudy (anonymous):

That will give you your answer in terms of pi

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Yes, so our answer is \(9 \pi mm^2\)

OpenStudy (anonymous):

yes pi3^3 = pi9

OpenStudy (anonymous):

wait would it be A

OpenStudy (anonymous):

wait its b? @iGreen.